ARC2580 STRUCTURES 1 HOMEWORK DUE 7-1-20 S. BRASGALLA, INSTRUCTOR NAME 18- 18-0 Girder G-4 20-0 BEAM B-2 JOIST J-1 Typic
Answer #1

Solution:

No Given :- Length of Joist span= 18 ft. spacing beth adjecent joist=14- 1.333 ft. Beam B3= span - 20 ft. D. L=250sf .. L.LMiho. Beam B3 : • No. of Joist on B3 = 20 ft 1.333 ft -1-15 Nos. 1 = 14 Nos Joisty No-14; P=7206b. -15@1.3377 L=2oft @ Vyman